VALIDATING WAVEFUNCTIONS IN MOMENTUM SPACE: A COMPUTATIONAL AND EXPERIMENTAL ELECTRON MOMENTUM SPECTROSCOPIC STUDY OF 2,6-STELLADIONE

\underline{FENG WANG}, Centre for Molecular Simulation and School of Information Technology, Swinburne University of Technology, PO Box 218, Melbourne, Victoria 3122, Australia; MICHAEL J. BRUNGER, KATE L. NIXON, LAURENCR CAMPBELL, School of Chemistry, Physics and Earth Sciences, The Flinders University of South Australia, GPO Box 2100, Adelaide, SA 5001, Australia.; DAVE A.WINKLER, Division of Molecular Sciences, CSIRO, Private Bag 10, Clayton South MDC, Victoria 3169, Australia.; BILL APPELBE, Victorian Partnership for Advanced Computing, PO Box 201, Carlton South, Vic. 3053, Australia; and ROLF GLEITER, Organic Chemistry Institute, University of Heidelberg, Inn Neunheimer Feld 270, D-69120, Germany.

The ground electronic state for 2,6-stelladione (C_{8}H_{8}O_{2}, X^{1}A_{1}), has been studied using quantum mechanical calculations such as density functional theory (DFT) calculations. Based on these calculations and the plane wave impluse approximation (PWIA), the momentum distributions (MDs) of a set of representable outer valence MOs for the 2,6-stelladione have been simulated. The simulated MDs are compared against our first time measurements using high-resolution electron momentum spectroscopy (HREMS). Such a detailed comparison between the experimental MDs and simulated MDs provides insight and understanding of the electron correlation effects as well as the orbital relaxation effects on the MOs. In addition, the accuracy of the DFT exchange-correlation (XC) functionals used and the basis sets tested.